### v1

# Mitigating Negative Side Effects in Multi-Agent Systems Using Blame Assignment

Abstract
When agents that are independently trained (or designed) to complete their individual tasks are deployed in a shared environment, their joint actions may produce negative side effects (NSEs). As their training does not account for the behavior of other agents or their joint action effects on the environment, the agents have no prior knowledge of the NSEs of their actions. We model the problem of mitigating NSEs in a cooperative multi-agent system as a Lexicographic Decentralized Markov Decision Process with two objectives. The agents must optimize the completion of their assigned tasks while mitigating NSEs. We assume independence of transitions and rewards with respect to the agents’ tasks but the joint NSE penalty creates a form of dependence in this setting. To improve scalability, the joint NSE penalty is decomposed into individual penalties for each agent using credit assignment, which facilitates decentralized policy computation. Our results in simulation on three domains demonstrate the effectiveness and scalability of our approach in mitigating NSEs by updating the policies of a subset of agents in the system.
1 Introduction
Many real-world settings involve the operation of multiple autonomous agents that do not require coordination for the completion of their individual assigned tasks in an environment, such as robots with distinct areas of operation in factory warehouses [?]. Such agents are often trained in isolation and their training does not account for potential negative interactions with other agents or their cumulative effects on the environment. Consequently, the agents may act in ways that are collectively undesirable when deployed, such as producing negative side effects (NSEs) [??]. This paper focuses on mitigating NSEs in cooperative multi-agent settings where the agents produce no (or negligible) NSEs when executing their policy in isolation, but their joint policy results in NSEs.
Consider warehouse robots, operating alongside humans, that optimize moving shelves between two locations. Each robot’s model provides the necessary information, including reward and transition dynamics, to complete its task optimally. The models may lack information about the effects of agents’ joint actions in the environment, such as a narrow corridor being blocked for human access when multiple robots simultaneously move large shelves through it. Thus, even when the agents are adept at completing their tasks and produce no NSEs when acting in isolation, their joint actions may have undesirable effects.
Figure 1: Overview of our solution approach. The agents independently compute optimal policies to complete their assigned tasks described by $R_{1}^{i}$ (Naive policy). The NSE Monitor computes the NSE penalty associated with the joint policy $\vec{\pi}$. The Blame Resolver then performs blame assignment to decompose the penalty in each joint state-action pair into penalties associated with each agent’s local state-action pair, which yields $R^{i}_{N}$. The agents finally recompute their policies by solving the bi-objective problem using lexicographic value iteration (LVI) with $R_{1}^{i}\succ R_{N}^{i}$.
Mitigating NSEs in multi-agent settings is challenging because: (1) NSEs are often discovered after deployment, since they depend on the environment, agent interactions and their assigned tasks; (2) NSEs and corresponding penalties are defined over joint actions, introducing a dependency between agents; and (3) the computational complexity of mitigating NSEs, without significantly affecting task completion, increases with the number of agents in the system. Prior research on NSEs primarily target single-agent settings [????]. It is not straightforward to apply them to multi-agent settings as they do not account for the agent interactions that produce NSEs. Recently, [?] proposed an approach to avoid NSEs by optimizing the future returns by other agents in the environment. However, it cannot avoid NSEs that require policy update for multiple agents and do not fully scale to large settings.
We formulate the problem of mitigating NSEs in a multi-agent system as a decentralized Markov decision process (DEC-MDP) with two objectives and a lexicographic ordering over them. The primary objective for each agent is to optimize its assigned task and the secondary objective is to minimize the NSEs. We assume independence of transition and rewards with respect to the agents’ assigned tasks [?]. Each agent’s reward for task completion is determined by its local state and actions and the overall reward for the system is the sum of individual agent’s rewards, and agents’ state transitions are affected only by their own actions. We use a lexicographic formulation since it is more intuitive and many real-world problems have an inherent ordering over objectives. It has been also been shown that lexicographic MDPs can be converted to a constrained MDP [?].
The agents do not have prior knowledge about NSEs but incur a joint penalty for them when executing certain actions together. Thus, while agents can independently compute policies to complete their tasks, mitigating NSEs requires addressing a form of dependence created by the joint penalty.
We present a metareasoning [??] approach to detect and mitigate NSEs. The metareasoner is a centralized entity that monitors the agents’ behaviors and has two components: (1) NSE Monitor that estimates the NSEs associated with agents’ joint policy, and outputs the corresponding penalty, and (2) a Blame Resolver that decomposes the joint penalty into individual agent penalties, using our algorithm Reward Estimation using Counterfactual Neighbors (RECON). It is assumed that the NSE monitor has access to a model of NSEs and the associated penalty, either provided as part of its design or acquired through human feedback. The blame resolver, using RECON, performs blame (credit) assignment to determine each agent’s relative contribution towards the NSE, based on which the joint NSE penalty is decomposed into local penalties for each agent [???]. This decomposition facilitates updating the model of each agent with NSE information via a penalty function, thereby enabling decentralized policy computation to mitigate NSEs.
Our solution framework uses a four-step approach to mitigate NSEs (Figure 1): (1) the agents first calculate optimal policies to complete their assigned tasks (referred to as naive policies); (2) the NSE monitor estimates the NSE penalty associated with the joint policy of the agents; (3) the blame resolver then decomposes the joint penalty into individual penalties for each agent, using blame (credit) assignment; and (4) the agents recompute their policies by solving a decentralized, bi-objective problem, with the prescribed reward for their task and the estimated local penalty for NSE, using lexicographic value iteration [?].
Our primary contributions are: (1) formalizing the problem of mitigating NSE in cooperative multi-agent settings as a decentralized, bi-objective problem; (2) presenting a metareasoning approach to update agent models with a penalty function corresponding to NSEs; (3) introducing an algorithm to decompose joint NSE penalty into individual penalties, using counterfactual-based credit assignment; and (4) empirical evaluation on three domains in simulation.
Credit Assignment Technique    Scalable Supports decentralized Compatible with Can generate NSE-
planning heterogeneous agents specific counterfactuals
Shapley Value
Difference Reward
Action Not Taken
D++
Wonderful Life Utility
Aristrocratic Utility
Value Decomposition
Our Approach
Table 1: Overview of different credit assignment techniques for mitigating side effects. A technique is scalable if it is computationally inexpensive to generate counterfactuals for problems with a few hundreds of agents. “--” indicates that the approach cannot be applied directly but can be modified to meet the requirements of our setting.
2 Related Works
Negative Side Effects
Prior works have considered different definitions of NSEs such as hindering the completion of future tasks or the operation of other agents [????], and model incompleteness in the form of missing information or incomplete reward function [????]. Our setting is similar to the latter works in that the agents have incomplete information about the cumulative effects of their joint actions. A common approach in such settings is to modify the existing reward function or to consider a secondary reward function associated with NSEs [????] and solve a multi-objective problem. The information about NSEs may be gathered using human feedback [??] . However, these approaches have been presented for a single agent setting or consider other agents as part of the environment, and it is not straightforward to extend them to multi-agent settings where multiple agent policies must be updated. Recently, a distributed coordination approach has been proposed for mitigating NSEs in multi-agent settings [?]. They target NSEs arising due to local agent interactions and formulate it as a distributed constraint optimization problem and the agents learn a policy to avoid NSEs using coordinated Q-learning. Their approach is not scalable as distributed constraint optimization can become intractable for large problems [?]. We present a metareasoning approach that scales to large multi-agent systems.
Credit Assignment.
Credit assignment is a popular approach to measure the contribution of an agent to team performance, which is used to convert a joint reward into individual agent rewards [?]. Difference Reward [?] and its variants such as D++ [?], Wonderful Life Utility and Aristrocratic Utility [?], perform credit assignment by comparing the joint rewards before and after the agent is removed from the system. Another type of credit assignment uses Shapley value and involves computing a value function for each agent by considering all combinations of possible agent interactions [????]. Table 1 summarizes the characteristics of different credit assignment techniques. Blame assignment for NSEs must be based on counterfactuals generated by considering NSE-relevant features only. The existing methods calculate counterfactuals by considering all state features which leads to incorrect blame assignment for NSEs.
3 Preliminaries
Decentralized Markov Decision Process (Dec-MDP).
A Dec-MDP is widely used to model decentralized multi-agent decision-making problems [?]. A Dec-MDP is defined by the tuple $\langle\mathcal{A},S,A,T,R\rangle$ with $\mathcal{A}$ denoting the finite set of $k$ agents in the system; $S\!=\!\hat{S}_{1}\!\times...\times\!\hat{S}_{k}$ denoting the joint state space, where $\hat{S}_{i}$ denotes the state space of agent $i$; $A=\hat{A}_{1}\!\times\ldots\!\times\!\hat{A}_{k}$ denoting the joint action space, where $\hat{A}_{i}$ denotes agent $i$’s actions; $T:S\times A\times S\rightarrow[0,1]$ denoting the transition function; and $R$ denoting the reward function. A joint policy $\vec{\pi}\!=\!(\pi_{1},...,\pi_{k})$ is a set of policies, one for each agent in the system.
Transition and reward independence.   A Dec-MDP with transition independence and reward independence [?] is a class of problems in which agents operate independently but are tied together through a reward structure that depends on all of their execution histories. A transition and reward-independent MDP satisfies the following:
$T(\hat{s}_{i}^{\prime}|{s},{a},{s}^{\prime}_{-i})=T_{i}(\hat{s}_{i}^{\prime}|\hat{s}_{i},\hat{a}_{i}),\forall i\in\mathcal{A}\vspace{-3pt}$
$R({s},{a},{s}^{\prime})=\sum_{i\in\mathcal{A}}R_{i}(\hat{s}_{i},\hat{a}_{i},\hat{s}_{i}^{\prime}).$
A transition and reward-independent Dec-MDP can be solved as $k$ single agent MDPs [?].
Lexicographic MDP (LMDP).
LMDPs are particularly convenient to model problems with potentially competing objectives with an inherent lexicographic ordering over them, such as ours where task completion is prioritized over NSE mitigation [??]. An LMDP is denoted by the tuple $M=\langle S,A,T,\boldsymbol{\mathrm{R}},o\rangle$ with finite set of states $S$, finite set of actions $A$, transition function denoted by $T\!:\!S\!\times\!A\times\!S\!\rightarrow\![0,1]$ and a vector of reward functions $\boldsymbol{\mathrm{R}}\!=\![R_{1},...,R_{k}]^{T}$ with $R_{i}\!:\!S\!\times\!A\!\rightarrow\!\mathbb{R}$, and $o$ denotes the strict preference ordering over the $k$ objectives. The set of value functions is denoted by $\boldsymbol{V}\!=\![V_{1},...,V_{k}]^{T}$, with $V_{i}$ corresponding to $o_{i}$,
$\boldsymbol{V}^{\pi}(s)\!=\!\boldsymbol{\mathrm{R}}(s,\pi(s))+\gamma\sum_{s^{\prime}\in S}T(s,\pi(s),s^{\prime})\boldsymbol{V}^{\pi}(s^{\prime}),\forall s\in S.$
A slack $\boldsymbol{\Delta}\!=\!\langle\delta_{1},...,\delta_{k}\rangle$ with $\delta_{i}\!\geq\!0$, denotes the acceptable deviation from the optimal expected reward for objective $o_{i}$ so as to improve the lower priority objectives. Objectives are processed in the lexicographic order. The set of restricted actions for $o_{i+1}$ is $A_{i+1}(s)\!=\!\{a\!\in\!A|\max\limits_{a^{\prime}\in A_{i}}Q_{i}(s,a^{\prime})-Q_{i}(s,a)\leq\eta_{i}$ where $\eta_{i}\!=\!(1-\gamma)\delta_{i}$,$\gamma\in[0,1)$. Refer [?] for a detailed background on LMDP.
4 Problem Formulation
Problem Setting.  Consider a setting where $m$ agents are independently performing their assigned tasks which is their primary objective $o_{1}\!=\!\{o_{1}^{1},...,o_{1}^{m}\}$. The agents operate based on a transition and reward independent decentralized Markov decision process (Dec-MDP), denoted by $M$ [?]. A meta-level process, metareasoner, monitors and controls the agents performance (object-level process), by intervening if agent actions have undesirable consequences [??].
While the model $M$ contains all the information necessary to optimize $o_{1}$, it may not describe the joint effects of agents’ actions, since it is irrelevant to task completion in this setting. Due to the limited fidelity of $M$, negative side effects (NSEs) occur when agents execute their policies jointly. The agents incur a penalty for NSEs, determined by a function $R_{N}$ which is known to the metareasoner (either by design or learned using human feedback) but unknown to the agents.
We make the following assumptions about the NSEs: (1) there are no (or negligible) NSEs when agents execute their policy in isolation but their joint policy produces NSEs that must be mitigated to the extent possible; (2) agents have no prior knowledge about NSEs of the joint actions, besides receiving a penalty assigned by the metareasoner; and (3) NSEs do not interfere with the agents’ task completion—they are undesirable but are not catastrophic.We target settings where the completion of the agents’ assigned tasks ($o_{1}$) are prioritized over minimizing NSEs ($o_{2}$), $o_{1}\!\succ\!o_{2}$.
MASE-MDP.  The problem of mitigating NSEs in a cooperative multi-agent system is formulated as a Dec-MDP with two objectives and a lexicographic ordering over them.
Definition 1.
A multi-agent side effects MDP (MASE-MDP) is a bi-objective Dec-MDP with lexicographic ordering over the objectives, denoted by $M=\langle\mathcal{A},S,A,T,\bm{R},o\rangle$, where:
  • $\mathcal{A}=\{1,\ldots,m\}$ is a finite set of agents in the system;
  • $S=\hat{S}_{1}\times\ldots\times\hat{S}_{m}$ is the joint state space;
  • $A=\hat{A}_{1}\times\ldots\times\hat{A}_{m}$ is the joint action space;
  • $T:S\times A\times S\rightarrow[0,1]$ is the transition function;
  • $\bm{R}\!=\![R_{1},R_{N}]$ is the reward function with $R_{1}\!:\!S\!\times\!A\!\rightarrow\!\mathbb{R}$ denoting the reward for task performance and $R_{N}\!:\!S\!\times\!A\!\rightarrow\!\mathbb{R}$ denoting the penalty function for NSEs; and
  • $o=[o_{1},o_{2}]$ denotes the objectives, where $o_{1}$ is the primary objective denoting agents’ assigned tasks and $o_{2}$ is minimizing NSEs, with $o_{1}\succ o_{2}$.
We consider independence of transition function and task completion reward $R_{1}$ in the MASE-MDP, meaning each agent’s transitions and task reward depend only on its local state and action. Hence, each agent can independently compute a policy for its assigned task by solving an MDPs [??]. However, the agents incur a joint penalty for the NSEs, which introduces a form of dependence and prevents decentralized computation of individual agent policies. In Section 5, we present an approach to decompose joint penalties into individual penalties, thereby facilitating decentralized planning.
Local and Global State Features.  We consider a factored state representation. Let $F$ denote the set of features in the environment, which are categorized into local features $F_{l}$ and global features $F_{g}$, $F\!=\!F_{l}\cup\!F_{g}$. The local features of an agent $i$ are denoted by $F^{i}_{l}$. Local features are agent-specific features that are controlled by the agent’s actions and affect its performance (e.g. an agent’s $x,y$ location). Global features are shared among agents and denote the overall state of the system. They are further divided into static and dynamic global features, denoted by $F_{gs}$ and $F_{gd}$ respectively, based on whether they are exogenous or can be modified by agent actions. An agent’s state $\hat{s}$ is described by $\vec{f}\!=\vec{f}_{l}^{i}\cup\vec{f}_{gd}\cup\vec{f}_{gs}$.
Definition 2.
Static global features $F_{gs}$ are exogenous factors that affect all agents and are observable to all agents but not changed by the agents’ actions (e.g. ocean currents).
Definition 3.
Dynamic global features $F_{gd}$ describe properties of the environment that affect agent operation directly or indirectly, and can be modified by agent actions (e.g. locations and sizes of the shelves moved by the agents).
In this work, the penalty for NSE associated with a joint action is modeled as a function of the dynamic global features of the joint state, $R_{N}({s})=\Omega(\vec{f}_{gd})$. However, the proposed framework and approach can be extended to handle penalties defined over joint states and actions.
5 Penalty Decomposition via Blame Assignment
Our algorithm for NSE penalty decomposition, Reward Estimation using Counterfactual Neighbors (RECON), is outlined in Algorithm 1. RECON first initializes each agent’s penalty function $R_{N}^{i}$ to zero (Line 1). After the agents have calculated their naive policies independently, the NSE Monitor component in the metareasoner calculates the joint penalty $\vec{r}_{N}$ for NSEs incurred from joint policy $\vec{\pi}$ (Line 2). If the penalty exceeds a pre-defined NSE tolerance threshold $\eta$, then the Blame Resolver decomposes the joint penalty into local penalties for each agent, based on their relative contribution to NSE, calculated using blame (credit) assignment (Lines 3-5). The penalty decomposition resolves the dependency induced by the joint penalty and enables decentralized policy computation to optimize task completion while minimizing NSEs.
Blame Estimation. Blame (credit) assignment estimates an agent’s contribution to a global penalty (or reward) using counterfactual states. The blame $B_{i}$ (Eqn. 1) is calculated for each agent $i$ in the joint state $s$, based on the difference between the current NSE penalty and the minimum NSE penalty that could have been achieved by the agent. The existing credit assignment techniques calculate counterfactual states by considering all state features. In our setting, NSE occurrence is determined only by the dynamic global features $(\vec{f}_{gd})$. Therefore, the counterfactuals must be calculated only over $\vec{f}_{gd}$. Generating counterfactual states for NSEs by considering all the state features could result in incorrect attribution. Consider multiple warehouse robots, with some carrying large shelves, navigating in a narrow corridor. The NSE of blocked paths for human access is determined by number of agents carrying large shelves. Generating counterfactual states traditionally will involve changing agent location but that does not provide information on NSE associated with multiple agents carrying large shelves. Hence, counterfactuals must be generated over NSE-relevant state features. We present a modified formulation of credit assignment that uses counterfactual neighbors instead of counterfactual states.
Definition 4.
The counterfactual neighbors of a joint state $s$, denoted by $s_{c}$, is the set of all states that vary only in the values of dynamic global features $(\vec{f}_{gd})$ while all other feature values are same as in $s$.
Counterfactual neighbors are therefore a subset of counterfactual states. Since the NSE penalty depends only on the state in our formulation, we calculate counterfactual neighbors over the state variables. The set of counterfactual neighbors that are reachable from the start state in the environment are referred to as valid counterfactual neighbors and are denoted by ${s}_{c}^{v}$.
Definition 5.
Agent-specific counterfactual neighbors, denoted by $s^{i}_{c}\!\subset s_{c}$, are states that differ in those dynamic global features that can be controlled by agent $i$, while other feature values are same as in the current joint state $s$.
Agent-specific counterfactuals allow us to estimate an individual agent’s contribution to NSE. Note that this is a typical form of counterfactual state generation for credit assignment, in that the agent contribution is estimated by either removing it or replacing its actions, while the other agents’ behaviors are fixed [??].
Algorithm 1 RECON
Input: NSE Tolerance $\eta$, $\vec{\pi}$ for primary assigned task
1:  Initialize $R^{i}_{N}(\hat{s})=0,\forall\hat{s}\in\hat{S}_{i},\forall i\in\mathcal{A}$
2:  Calculate joint penalty $r_{N}^{\vec{\pi}}$ for $\vec{\pi}$
3:  if $r_{N}^{\vec{\pi}}$ $>$ $\eta$ then
4:     Compute blame $B_{i}(s)$ using Eqn. 1, $\forall i\in\mathcal{A},\forall s\in S$
5:     $R^{i}_{N}(\hat{s}_{i})\leftarrow B_{i}(s)$, $\forall i\in\mathcal{A},\forall\hat{s}_{i}\in s,\forall s\in S$
6:  end if
7:  return   $[R^{1}_{N},...,R^{m}_{N}]$
Counterfactual neighbors vs state counterfactuals.
Let ${s}^{i,v}_{c}\subset s^{v}_{c}$ be the set of valid (reachable) counterfactual neighbors for agent $i$. The blame $B_{i}$ for agent $i$ in the joint state $s$ is calculated as,
$B_{i}({s})=\frac{b_{i}({s})}{\sum_{i\in\mathcal{A}}b_{i}(s)}\cdot{R_{N}}(s),\text{ with}$ (1)
$b_{i}(s)=\frac{1}{2}\left({R_{N}^{*}}+\epsilon+\left(R_{N}(s)-\min_{s^{\prime}\in s^{i,v}_{c}}R_{N}(s^{\prime})\right)\right),$ (2)
where $R_{N}$ is the joint NSE penalty function, $R_{N}^{*}=\max_{s\in S}R_{N}(s)$ denotes the maximum joint NSE penalty possible which is used to rescale blame in the range $[0,R_{N}^{*}]$ to avoid impractical values that might be negative or exceed the joint NSE penalty itself. $\epsilon$ is a small fixed value to avoid singularities during normalization and rescaling, and $b_{i}(s)$ is an intermediate value used to calculate $B_{i}(s)$. The above equation assigns blame proportional to the agent’s ability to mitigate NSEs. This ensures that agents already making their best efforts are penalized less compared to other agents.
The metareasoner’s Blame Resolver compiles a local penalty function $R^{i}_{N}$ for each agent $i$ as follows:
$R^{i}_{N}(\hat{s}_{i})=B_{i}(s),\forall\hat{s}_{i}\in s,\forall s\in S,i\in\mathcal{A}.$ (3)
The agents then solve the MASE-MDP, with the prescribed $R_{1}^{i}$ corresponding to their assigned task and $R^{i}_{N}$ provided by the metareasoner, using lexicographic value iteration (LVI) [?].
Generalizing $R^{i}_{N}$
The penalty function $R^{i}_{N}$ obtained from RECON in Algorithm 1, is calculated based on the blame values corresponding to $\vec{\pi}$ and does not provide any information about potential NSEs that may occur if the agents followed a different joint policy. As a result, NSEs may persist when the agents update their policy by solving the MASE-MDP with $R^{i}_{N}$. To overcome this limitation, we consider a supervised learning approach to generalize the NSE penalty to unseen situations by using the $R^{i}_{N}$, based on the initial $\vec{\pi}$, as training data. The prediction accuracy can further be improved by including the counterfactuals as part of the training [?]. In our experiments, learning to generalize using counterfactuals improves the performance in some settings, compared to generalizing without using counterfactual information. Generalization helps scale the algorithm to systems with large number of agents and eliminates the requirement of having an iterative approach where agents could discover new NSE in unseen states. An iterative RECON will work but it incurs a high computational overhead and does not guarantee a monotonic improvement in policy with each iteration.
(a) Salp agents Refer to caption, inspired from [??], are tasked with collecting physical samples from sites Refer to caption,Refer to caption, and depositing them at lab facility Refer to caption surrounded by corals Refer to caption prone to damage.
(b) Overcooked environment, inspired from [?], shows 4 agents tasked with soup orders. The kitchen shown above has tomatoes Refer to caption, onions Refer to caption, cooking pots Refer to caption, clean dishes Refer to caption, and a serving counter Refer to caption, dirty dishes Refer to caption, and garbage bins Refer to caption for waste.
(c) Warehouse environment, inspired from [?], shows agents Refer to caption tasked with processing shelves Refer to caption, at counter Refer to caption. The narrow corridors Refer to caption with human workers Refer to caption that are inconvenienced around multiple robots carrying shelves.
Figure 2: Example problem instances from domains used in our experiments.
6 Experimental Setup
We evaluate our approach in simulation using three proof-of-concept grid-world domains (salp, overcooked, and warehouse). Figure 2 shows sample instances from each domain. All algorithms are implemented in Python and the simulations were run on an Ubuntu machine with $32$ GB RAM.The MASE-MDP problem is solved using lexicographic value iteration [?] and results are averaged over five instances in each domain. We test with zero NSE tolerance $\eta\!=\!0$ to evaluate the extent of NSE mitigation using RECON, and using $\epsilon=10^{-4}$ for rescaling the blame values.
Baselines
We compare the performance of RECON with four baselines. First is a Naive policy that is optimal for agent task completion and does not optimize NSE mitigation, providing an upper bound on NSE penalty. Second is performing credit assignment using the Difference Reward technique [?] that calculates blame by removing the agent from the system, $B_{i}(s)=R_{N}(s)-R_{N}(s_{-i})$. A direct comparison with this technique is not feasible since it calculates counterfactuals based on assumptions that do not hold in our setting. Therefore, we modify the approach to calculate counterfactuals over dynamic global features of the agents, and calculate blame as $B_{i}(s)=R_{N}(s)-\max_{s^{\prime}\in s^{i,v}_{c}}R_{N}(s^{\prime})$. We use max over all valid counterfactual neighbors to get an aggressive blame assignment, producing conservative (safer) behavior. Third, we compare with a variant of considerate reward inspired from  [?], where they augment agents’ reward functions with value functions of subsequently acting agent to avoid their definition of NSE. Since in our setting, agents operate simultaneously and NSE is defined for multiple agents, we augment agents’ prescribed rewards with blame contributions from the rest of the agents as shown in Equation 4.
$R^{i}_{r}(\hat{s}_{i})=\alpha_{1}\frac{R_{1}^{i}(\hat{s}_{i})}{R_{1}^{*}}+\alpha_{2}\frac{R_{N}(s)-B_{i}(s)}{R_{N}^{*}}$ (4)
where $R_{1}^{*}=\max_{\hat{s}\in\hat{S}_{i}}R_{1}(\hat{s})$ is the maximum possible reward for assigned task for agent $i$, $\alpha_{1}$ and $\alpha_{2}$ are the selfish and care coefficients respectively. $R_{1}^{*}$ and $R_{N}^{*}$ are used to normalization so that their relative scales are not an inherent factor, but a controlled one (using $\alpha_{1}$, $\alpha_{2}$). Fourth, we compare with generalized RECON w/o counterfactual data, which uses only the $R^{i}_{N}$ values from RECON to generalize the penalty situations using supervised learning on global features of the local states and attributed blame values. Finally, we compare with generalized RECON w/ counterfactual (cf) data, which uses both counterfactuals and $R^{i}_{N}$ values to generalize $R^{i}_{N}$ to unseen situations using supervised learning similar to the previous baseline. The techniques are evaluated on the following domains.
Sample Collection using Salps
Salps are marine organisms that can form chains and coordinate movements [?]. We consider a setting where salp-inspired underwater robots must collect chemical samples from different locations in the seabed. Each agent is tasked with collecting a specific type of sample and depositing it to a designated location for analysis. An agent’s state is denoted by $\langle x,y,sample,coral,status\rangle$, where $x,y$ denote agent’s location, $sample$ indicates the sample type, $coral$ indicates presence of coral at $x,y$, and $status$ is a binary variable indicating if the sample has been deposited at the destination. In our experiments, samples can be of type $A,B$ or $X$ indicating no sample. Agents can move in all four directions, and $pick$ and $drop$ samples, each with a reward of $-1$. The reward for task completion is $+100$. The move actions succeed with a probability of $0.8$ or fail and slide to an adjacent location with a probability of $0.2$. Features used for training when generalizing $R^{i}_{N}$ are $\langle sample,coral\rangle$. We test with five instances of $20\!\times\!20$ grids that vary in coral locations. This allows us to evaluate the different approaches to mitigate NSEs, while keeping other conditions fixed.
NSE: The salp-like robots may have chemical residues floating around them when transporting samples. A joint NSE occurs when multiple robots carrying chemical samples are in immediate vicinity of corals, potentially damaging it.
Overcooked
In this domain, agents are tasked with preparing and serving food [?]. The agents work together to deliver a fixed number of tomato and onion soup orders, while keeping the kitchen clean. Each agent is assigned a cleaning or a cooking task, and it receives a reward of $+100$ upon task completion. Actions in non-goal state have a reward of $-1$. In each problem instance, $20\%$ agents are assigned to cleaning and the rest are assigned the cooking task. An agent’s state is represented as $\langle x,y,dir,object,bin,status\rangle$, where $x,y$ denote its location, $dir$ denotes its orientation, $bin$ indicates the presence of garbage bins at $x,y$, and $status$ indicates the completion status of its assigned task. Agents can move forward in all directions, and $interact$ with objects in the kitchen. Interactions vary from picking and putting down ingredients and dishes, to using the cooking pots and dumping garbage in bins. Agent’s move actions succeed with a probability of $0.8$ or slide to an adjacent location with $0.2$ probability. The interact actions succeed with probability of $0.8$ or fail and remain in the same state with probability $0.2$. Features used for generalizing $R^{i}_{N}$ are $\langle object,bin\rangle$. We test with five instances of $15\!\times 15$ grids that vary in the locations of the garbage bins. NSE: The garbage bins emit bad odors and attract flies. Any object involved in soup preparation must therefore be kept away from the garbage bins or the waste must be disposed at a farther bin, otherwise the food quality is affected.
(a) Salp domain
(b) Overcooked domain
(c) Warehouse domain
Figure 3: Average NSE penalty and standard deviation for varying % of agents undergoing policy update in each domain with $25$ agents.
Warehouse Inventory Management
[…]
NSE: When multiple robots carrying shelves are simultaneously in the same narrow corridor, it inconveniences human workers to navigate and access the area. The NSE penalty depends on shelf size and the number of agents carrying shelves in the same corridor.
(a) Salp domain
(b) Overcooked domain
(c) Warehouse domain
Figure 4: Average NSE penalties and standard deviations, averaged over five problem instances and varying number of agents.
(a) Salp domain
(b) Overcooked domain
(c) Warehouse domain
Figure 5: Average runtime of all techniques, with increasing number of agents across five problem instances, with standard deviation.
NSE Penalty
We employ a logarithmic NSE penalty function (Equation 5) for the domains described earlier. The NSE penalty calculated as,
$R_{N}(s)=\sum_{d\in F_{gd}}\sum_{k\in d}\beta_{k}\cdot\log\left(\alpha_{d}N_{k}+1\right),$ (5)
where $F_{gd}$ is the set of dynamic global features in the environment, $d$ is a feature in $F_{gd}$, and $k$ is the value of feature $d$. $N_{k}$ is the number of agents with dynamic global feature value $d\!=\!k$ in the joint state $s$. $\beta_{k}$ is a scaling factor that is directly related to the NSE penalty for a specific feature value $d=k$. Essentially, higher NSE penalties correspond to higher values of $\beta_{k}$, reflecting that mitigating more critical features yields a more substantial reduction in NSE penalty, aligning with the nuanced dynamics of real-world environments. $\alpha_{d}\,(>0)$ is a sensitivity parameter for NSE associated with feature $d$. A larger $\alpha_{d}$ denotes that the NSE penalty is more sensitive to an increasing number of agents with a specific feature value.
We use a logarithmic function to effectively model scenarios where the NSE impact plateaus with a certain number of agents performing undesirable actions [??]. For example, there may be no significant difference in the NSE impact caused by 10 agents blocking a corridor versus 11 agents blocking a corridor. A log penalty will lead to prioritizing the mitigation of severe NSEs, by updating the policy of a subset of agents. However, note that RECON can work with any penalty function in practice.
7 Results and Discussion
We compare the effectiveness of different techniques in mitigating NSEs, along multiple dimensions such as the number of agents undergoing policy update to mitigate the impact produced, NSE avoidance by generalizing $R^{i}_{N}$ to unseen situations, and run time of the techniques.
Effect of number of agents undergoing policy update     First, we evaluate the effectiveness of our approach in mitigating NSEs, by varying the number of agents whose policy must be updated after $R^{i}_{N}$ is calculated. For each domain, we consider $25$ agents in the environment and vary the percentage of agents undergoing policy update to minimize NSEs, from $10\%$ to $100\%$. We select agents to update policies by ranking them in the decreasing order of their blame values. The results, in Figure 3, show naive policy as a straight line in all the domains, since it is indifferent to NSEs. For the salp sample collection domain, Figure 3(a) shows that the NSE penalty plateaus at the $50\%$ mark, where the difference reward baseline shows its best performance against the naive policy. For the warehouse domain, in Figure 3(c), we see a similar trend of NSE penalty plateauing at the $50\%$ mark however, difference reward fails to mitigate NSE and performs same as the naive policy. In case of the overcooked domain, Figure 3(b) shows a steady decline of NSE penalty using RECON and its generalized versions but similar to the warehouse domain results, difference reward fails to mitigate NSE and performs same as the naive policy.
Note that in some cases, NSE may not be avoided even when we update the policies of $100\%$ of the agents in the system. This is because we prioritize completing the task optimally over minimizing NSEs. Hence in some cases, it may be impossible for the agents to avoid NSEs, while optimally completing their tasks. This is a problem characteristic and not a limitation of RECON.
Overall, the results show that RECON is able to mitigate NSEs, without updating the policies of a large number of agents. Based on these results, we update the policies of $50\%$ of agents in all the following experiments, across all domains, since we consistently see a plateau and drop in the NSE penalty at $50\%$. For some settings, updating all the agents’ policies may be effective but it is not practically feasible.
Effect of generalizing $R^{i}_{N}$    Since the penalty function $R^{i}_{N}$ is calculated based on the blame values corresponding to $\vec{\pi}$, policy updates may produce novel NSEs that were not encountered earlier. We evaluate the effectiveness of generalizing $R^{i}_{N}$ to unseen situations by comparing the performances of RECON, with and without generalization using counterfactual data, against the baselines. Based on results from Figure 3, we conduct experiments with $50\%$ of agents undergoing policy update for each technique. Figure 4 shows that both the generalized versions of RECON outperform other methods, consistently reducing NSE penalty across domains. In few instances, generalized RECON with counterfactual data outperforms the generalization without counterfactual data. These results show that generalization is useful and can mitigate NSEs considerably even when counterfactual data is not used for training. We also tested the techniques when $100\%$ of agents undergo policy update. We observe that the NSE penalties further reduce by over $60\%$. While this may be practically infeasible for large systems, it shows the effectiveness of our approach in mitigating NSEs.
Scalability     We evaluate the scalability of our approach in terms of run-time to solve the problems as we increase the number of agents in the system from $10$ to $100$. Figure 5 plots the runtime of different techniques, in the three domains, with varying number of agents. The run times increase in an approximately linear manner as the number of agents increases. The considerate reward baseline takes the least time possible because it involves solving only for one agent as other agents are treated as part of the environment. Difference reward baseline takes longer since policy computation is done in two stages but it does not evaluate all counterfactual neighbors and defaults to the counterfactual producing maximum NSE penalty. RECON takes longer than difference reward since it involves computing and comparing counterfactuals. As expected, RECON with $R_{N}$ generalizations take longer than the other approaches, since they generalize NSE penalties over the entire state space. However, the time increase is small, compared to the reduction in NSE penalties they offer, even in settings with $100$ agents. Since RECON is an offline approach, the slightly higher runtimes for additional processing to generalize $R_{N}$ may be tolerable in many settings. Since RECON is an offline approach, the slightly higher runtimes for additional processing to generalize $R^{i}_{N}$ may be tolerable in many settings.
8 Summary and Future Work
This paper formalizes the problem of mitigating NSEs in cooperative multi-agent settings as a decentralized, bi-objective problem. The agents’ assigned tasks follow transition and reward independence. The agents produce no NSE when operating in isolation but their joint actions produce NSE and incur a joint penalty. The agents have no prior knowledge about the side effects of their joint actions. We present a metareasoning approach to detect NSEs and update agent policies in a decentralized manner, by decomposing the joint NSE penalty into individual penalties. Our algorithm, RECON, uses a counterfactual-based blame attribution to estimate each agent’s contribution towards the joint penalty. Our experiments demonstrate the effectiveness of our approach in mitigating NSEs, by updating the policies of a subset of agents.
Our framework currently supports Dec-MDPs with transition and reward independence. In the future, we aim to relax this assumption and extend our approach to settings with tightly coupled task assignment. Another interesting future direction is to exploit the agent dependencies and interactions for task completion to mitigate NSEs, by leveraging their complementary skills.
Acknowledgments
This work was supported in part by ONR grant number N00014-23-1-2171.